# Bridge the Modality and Capacity Gaps in Vision-Language Model Selection

### v1

@@ 3.2 Bridge the Modality Gap and Capacity Gap
The modified text embedding $\tilde{\bm{t}}_{m,i}$ serves as better image proxies. In other words, classification metrics on $f^{T}_{m}(t_{i})$ only reveal the discerning ability of the text encoder of $f_{m}$, which is far from the (cross-modal) zero-shot classification ability due to the modality gap. By bridging such a gap with modified text embedding, classification metrics on $\tilde{\bm{t}}_{m,i}$ are closer to the classification metrics on images with textual classifier. Therefore, we use $\{\tilde{\bm{t}}_{m,1},\cdots\}$ in Equation 15 as better inputs to calculate $\bm{s}_{m,\mathcal{T}}$ in Equation 7. The updated score vectors $\bm{s}_{m,\mathcal{T}}$ are then input to the ranker model $f_{R}$, which is able to get more accurate VLM’s performance prediction $\hat{p}_{m,\mathcal{T}}$. Based on the performance prediction $\{\hat{p}_{m,\mathcal{T}}\}_{m=1}^{M}$, we can obtain the VLMs’ ranking:
$\bm{\hat{r}}_{k,\mathcal{T}}$ $=[\hat{r}_{1,\mathcal{T}},\cdots,\hat{r}_{M,\mathcal{T}}]$ (16)
$=\text{Ranking}\left(\left[\hat{p}_{1,\mathcal{T}},\cdots,p_{M,\mathcal{T}}\right]\right)\;.$
$\text{Ranking}(\cdot)$ transforms the accuracy values into their performance ranking within the VLM Zoo. $\bm{r}_{k}^{\mathcal{S}}\in\mathbb{Z}_{+}^{M}$ contains integer values from 1 to $M$.
@@
Bridge the Capacity Gap. Whether the $m$-th VLM $f_{m}$ fits the target task $\mathcal{T}$ could also be determined by the performance of $f_{m}$ on the open-source datasets related to $\mathcal{T}$. As we mentioned in subsection 2.3, the performance of a VLM varies from one dataset to another, so we calculate the VLM’s class-specific performance ranking over the whole open-source datasets. Given the $k$-th open-source class $c_{k}^{\mathcal{S}}$ and the corresponding set of images $D_{\mathcal{S}_{k}}^{I}$ as in Equation 12, we calculate the zero-shot classification accuracy $p_{m,k}^{\mathcal{S}}$ via Equation 2. Based on the performance rankings of all $M$ VLMs, we obtain each VLMs’ ranking on the $k$-th open-source class $c_{k}^{\mathcal{S}}$ using Equation 17. We calculate VLMs’ ranking for each of the $k_{\mathcal{S}}$ open-source classes similarly.
$\bm{r}_{k}^{\mathcal{S}}$ $=[r_{1,k}^{\mathcal{S}},\cdots,r_{M,k}^{\mathcal{S}}]$ (17)
$=\text{Ranking}\left(\left[p_{1,k}^{\mathcal{S}},\cdots,p_{M,k}^{\mathcal{S}}\right]\right)\;.$
@@
Next, we consider estimating the ranking of a certain VLM $f_{m}$ on the target dataset $\mathcal{T}$. By re-organizing the ranking values in Equation 17, the performance ranking vector of $f_{m}$ on all $k_{\mathcal{S}}$ open-source classes is $\bm{r}_{m}^{\mathcal{S}}=[r_{m,1}^{\mathcal{S}},\cdots,r_{m,k_{\mathcal{S}}}^{\mathcal{S}}]\in\mathbb{Z}_{+}^{k_{\mathcal{S}}}$. If $f_{m}$ has a higher ranking on certain open-source classes and those classes are related to the classes in the target set, the performance ranking of $f_{m}$ on the target dataset may also be high.
Thus, we perform a weighted sum of ranking values in $\bm{r}_{m}^{\mathcal{S}}$ and assign larger weights to those classes related to the target dataset. Since the semantic similarity between open-source and target classes is measured by $\bm{\gamma}^{*}$, we apply
$\hat{\bm{r}}_{m,\mathcal{T}}=\bm{r}_{m}^{\mathcal{S}}\;\bm{\gamma}^{*}\;.$ (18)
@@
Elements in $\hat{\bm{r}}_{m,\mathcal{T}}\in\mathbb{R}^{k_{\mathcal{T}}}$ are the predicted ranking of the m-th VLM $f_{m}$ for $k_{\mathcal{T}}$ target classes. Since we only need the relative order of ranks, there is no additional scale factor in Equation 18. After that, we average class-specific ranking values in $\hat{\bm{r}}_{m,\mathcal{T}}$, and use the mean of $\hat{\bm{r}}_{m,\mathcal{T}}$ to denote the ability of $f_{m}$ on the target dataset $\mathcal{T}$. In summary, we take account of the VLM ranks on related datasets, and such class-specific ranking estimation bridges the capacity gap in VLM selection.
@@ 3.3 Summary of Swab
As is described in section 3.2, given a target dataset $\mathcal{T}$ and a VLM $f_{m}$, we denote the two performance ranking predictions by bridging the modality gap and capacity gap as $\hat{r}^{1}_{m,\mathcal{T}}$ and $\hat{r}^{2}_{m,\mathcal{T}}$, respectively. $\hat{r}^{1}_{m,\mathcal{T}}$ is predicted based on ModelGPT with our modified embeddings of the generated auxiliary texts for $\mathcal{T}$. $\hat{r}^{2}_{m,\mathcal{T}}$ is predicted by the weighted sum of VLM’s class-specific ranking values on the open-source datasets. These two predictions respectively originate from non-learning-based and learning-based LOVM methods. We ensemble two predictions together and achieve a more accurate model ranking estimation. We utilize the weighted Borda Count to aggregate two rankings:
$\hat{r}^{\rm ens}_{m,\mathcal{T}}=\alpha\cdot\hat{r}^{1}_{m,\mathcal{T}}+\left(1-\alpha\right)\cdot\hat{r}^{2}_{m,\mathcal{T}}.$ (19)
Ultimately, Swab determines the final predicted ranking of the VLMs in the VLMs Zoo based on $\hat{r}^{\rm ens}_{m,\mathcal{T}}$. The pseudo-code of Swab is listed in Algorithm 1.
@@ 4 Experiments
We evaluate Swab on LOVM benchmark (Zohar et al., 2023) and analyze whether bridging the modality gap and capacity gap helps VLM selection.
@@ 4.1 Evaluation on LOVM Benchmark
Setups. We follow LOVM (Zohar et al., 2023) to use a VLM Zoo with 35 pre-trained VLMs, which differ in aspects such as model architecture, pre-training datasets and training methods. We evaluate different methods on 23 datasets, i.e. ImageNet (Deng et al., 2009), Aircraft (Maji et al., 2013), CIFAR100 (Krizhevsky & Hinton, 2009) and so on. We obtain VLM’s ground truth ranking based on VLM’s Top-1 Accuracy calculated on target dataset’s test image set.
@@
Baseline. We select representative methods for each of the three paradigms mentioned in Section 2.2 as our baselines. For the first paradigm, we use four classic model selection methods including H-Score (Bao et al., 2019), NCE (Tran et al., 2019), LEEP (Tran et al., 2020) and LogME (You et al., 2021). For the second paradigm, we use the ranking of VLM on ImageNet (INB) and the average ranking (Average Rank) on all datasets except the current target dataset in the LOVM Benchmark as our baselines, respectively. For the third paradigm, we compare our method with ModelGPT (Zohar et al., 2023).
@@
Evaluations. We use Top-5 Recall and Kendall’s Rank Correlation to measure the ranking similarity between the predicted model rankings and the ground truth ones to evaluate the VLM selection method’s performance. We also calculate the sum of these two metrics to consider the method’s comprehensive capability.
@@
Implementation Details. For a fair comparison, Swab follow ModelGPT (Zohar et al., 2023) to sequentially extract a target dataset from each of the 23 datasets in the LOVM Benchmark and treat the remaining datasets as open-source datasets. Besides, Swab adopts the approach outlined in ModelGPT to add Gaussian noise to corrupt the embeddings of the target dataset’s generated text data, which can improve the VLM selection method’s performance. Adding noise to features to boost the method’s performance has been used in prior work (Ye et al., 2017). For H-Score, NCE, LEEP, LogME, INB, and Average Rank, we follow the practices of previous work and do not add noise to the model’s inputs. To avoid the randomness introduced by the noise affecting the reliability of the experiment results, we conduct ten repeated experiments using random seeds from 1 to 10 and report the mean performance and standard deviation of different methods in Table 1.
@@
Table 1: Results on LOVM Benchmark. We evaluate our method’s performance over 23 datasets and 35 pre-trained VLMs. The results are averaged over all datasets. Our Swab achieves the best results across all metrics. For methods that involve adding random noise to data features, we report the standard deviation of metrics across 10 experiments to mitigate the impact of randomness on result reliability.
@@
H-Score 0.174 0.035 0.209
NCE 0.252 -0.029 0.223
LEEP 0.174 0.035 0.209
LogME 0.209 -0.029 0.180
INB 0.452 0.177 0.629
Average Rank 0.452 0.133 0.585
ModelGPT $\text{0.457}\scriptstyle\pm\text{0.009}$ $\text{0.203}\scriptstyle\pm\text{0.018}$ $\text{0.660}\scriptstyle\pm\text{0.015}$
Swab $\textbf{0.486}\scriptstyle\pm\text{0.010}$ $\textbf{0.268}\scriptstyle\pm\text{0.011}$ $\textbf{0.754}\scriptstyle\pm\text{0.003}$
Results Analysis. From Table 1, we can draw the following conclusions: (1) Metric-based non-learning model selection methods such as LogME show poor performance on the LOVM Benchmark. This is primarily because such algorithms rely on the target dataset’s images, thus the modality gap has a greater negative impact on them when using generated text data as a substitute for images. (2) Using open-source datasets is helpful for LOVM. We find that using open-source datasets in a non-learning way (e.g. INB, Average Rank) or a learning way (e.g. ModelGPT) all helps LOVM, since their performance significantly surpasses that of methods not utilizing open-source datasets (e.g. LogME). (3) Despite leveraging more open-source datasets, the performance of Average Rank is worse than INB. This confirms our analysis of the Capacity Gap, which suggests a discrepancy between the average ranking of a VLM and its ranking on a specific dataset. (4) Our Swab achieves the best performance across all evaluation metrics. Notably, our final performance of $R_{5}+\tau$ (0.754) represents a significant improvement of 9.4% over the original SoTA (State-of-The-Art) method ModelGPT (0.660).
@@ 4.2 Ablation Study
We conduct ablation studies to demonstrate that bridging the Modality Gap and Capacity Gap are both essential for Swab. Table 2 presents our experiment results, from which we can observe that Swab achieves the best performance when both gaps are bridged simultaneously. The ablation study confirms our analysis.
Table 2: Ablation Study of LOVM. Swab-C, Swab-M, and Swab indicates only bridging the Modality Gap, only bridging the Capacity Gap, and bridging both gaps in Swab.
@@
Swab-C $\text{0.478}\scriptstyle\pm\text{0.012}$ $\text{0.213}\scriptstyle\pm\text{0.011}$ $\text{0.691}\scriptstyle\pm\text{0.007}$
Swab-M $\text{0.459}\scriptstyle\pm\text{0.010}$ $\text{0.251}\scriptstyle\pm\text{0.001}$ $\text{0.710}\scriptstyle\pm\text{0.011}$
Swab $\textbf{0.486}\scriptstyle\pm\text{0.010}$ $\textbf{0.268}\scriptstyle\pm\text{0.011}$ $\textbf{0.754}\scriptstyle\pm\text{0.003}$
4.3 Influence of Key Components in Swab
Will Bridge the Capacity Gap Be Beneficial for VLM Selection? We compare the LOVM performance directly using the VLM’s average ranking on each class of open-source datasets and weighted-sum ranking based on transport matrix $\bm{\gamma}^{*}$. The results are shown in Table 3. From the results, we can find that utilizing class relevance to bridge the Capacity Gap is beneficial for VLM’s Model Selection.
Table 3: Compare results of $\hat{r}^{2}_{m,\mathcal{T}}$ on the LOVM Benchmark before and after bridging the Capacity Gap.
@@ 4.3 Influence of Key Components in Swab
Average Rank 0.452 0.133 0.585
OT Weighted Rank 0.443 0.275 0.718
Will Bridge the Modality Gap Be Beneficial for VLM Selection? To eliminate the interference of other factors, we solely utilize the learning-based predicted rankings $\hat{r}^{1}_{m,\mathcal{T}}$ in Swab, and the input to the ranker model $f_{m}$ consists only of metrics calculated on the generated text data $\tilde{D}_{\mathcal{T}}$. In this way, the performance of the method depends solely on the quality of the generated text data $\tilde{D}_{\mathcal{T}}$. From the Table 4, we can find that the generated text data $\tilde{D}_{\mathcal{T}}$ become better image proxies after bridging the Modality Gap.
Table 4: Compare results of $\hat{r}^{2}_{m,\mathcal{T}}$ on the LOVM Benchmark before and after bridging the Modality Gap (MG).
Method $R_{5}(\uparrow)$ $\tau(\uparrow)$ $R_{5}$ + $\tau(\uparrow)$
Before Bridging MG 0.249 0.064 0.313
After Bridging MG 0.350 0.196 0.546
Which Kind of Gap Vectors Should We Use? When utilizing the gap vectors from open-source datasets, we have two options: (1) Use the dataset-level mean gap vector calculated on the whole dataset’s image-text pairs. (2) Use the class-level mean gap vector calculated on the corresponding class’s image-text pairs. We hope that the gap vectors used to calculate the mean gap vector are as close to each other as possible, so that their mean vector can serve as a good substitute for the whole set. Based on this idea, we examine the statistical properties of dataset-level gap vectors set and class-level gap vectors set, respectively. We calculate three metrics which include: (1) the standard deviation of these gap vectors’ magnitude; (2) the mean cosine similarity between these gap vectors and their corresponding mean gap vectors; and (3) the standard deviation of these cosine similarities. These metrics reflect the consistency of the gap vectors. Table 5 shows the results.
Table 5: Results of metrics measuring gap vectors’ consistency. M: Magnitude, D: Direction. Metrics are averaged across datasets and classes, respectively.
@@
Class Mean 0.03 0.85 0.05
From the Table 5, we can find that the class-level gap vectors tend to be more consistent, which inspires us to use the class-level mean gap vectors. We also compare the results of $\{\hat{r}^{1}_{m,\mathcal{T}}\}_{m=1}^{M}$ on LOVM Benchmark using the dataset-level mean gap vectors and the class-level mean gap vectors, respectively. The implementation details are the same as Table 4. Table 6 shows the results, which verifies that using the class-level mean gap vectors is a better choice.
Table 6: Results of $\hat{r}^{1}_{m,\mathcal{T}}$ on the LOVM Benchmark using the dataset-level mean gap vectors and class-level mean gap vectors.
@@
Dataset Mean 0.338 0.091 0.429
Class Mean 0.350 0.196 0.546
Figure 4: The Spider Charts of a VLM. The Spider Chart measures the model’s capabilities across different domains. The axis represents the model’s average ranking on datasets in different domains. The greater the overlap between the predicted VLM’s spider chart and the actual VLM’s spider chart, the better the model selection method is. Our Swab performs better than ModelGPT.
4.4 Visualizing Swab’s Results by Spider Chart
We select some datasets from six different domains out of the 23 datasets in LOVM Benchmark (see Appendix C.4 for more details). We take the average of the model’s rankings across multiple datasets within the same domain as the model’s ranking in this domain. Afterward, we use ModelGPT and Swab, respectively, to predict the VLM rankings on datasets across different domains, thereby obtaining the predicted rankings of the VLM in six domains. We use a spider chart to visualize the gap between the VLM’s predicted rankings and actual rankings. The different vertices of the spider chart represent the VLM’s ranking in different domains. The score at each vertex indicates the model’s specific ranking. The higher the ranking (closer to 1), the better the model’s performance in that category. Figure 4 displays the spider chart for OpenAI’s CLIP RN50×64. By comparing the VLM’s predicted rankings generated by Swab and ModelGPT with the actual VLM rankings, we find that Swab is more capable of accurately estimating a VLM’s abilities compared to ModelGPT, and the abilities estimated by Swab are very close to the model’s true capabilities.
5 Related Work
Vision-Language Models. Vision-Language Models represent a class of multimodal models adept at correlating textual and visual information. VLMs are pre-trained on extensive text-image pairs using contrastive loss, endowing them with powerful text-image matching capability. Prominent VLMs include CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), FLAVA (Singh et al., 2022), Florence (Yuan et al., 2021), and GLIP (Li et al., 2022). VLMs possess robust zero-shot image classification capabilities (Radford et al., 2021), which enables its widespread application in tasks characterized by long-tail distributions or those where collecting substantial training data is challenging, such as medical image analysis. Some work (Menon & Vondrick, 2023; Yi et al., 2024) has also shown that by incorporating external knowledge, the zero-shot capabilities of CLIP can be further enhanced. In recent years, the number of open-source VLMs has been increasing (Ilharco et al., 2021). Previous work (Zohar et al., 2023) has pointed out that different VLMs possess varying image classification capabilities. This indicates that the performance of VLMs can vary significantly across different tasks and domains. These models with diverse capabilities constitute a VLM Zoo rich in knowledge. This VLM Zoo enables us to utilize different VLMs for various classification tasks, thereby changing the paradigm of using a single VLM to complete diverse classification tasks. This paper focuses on selecting the most suitable VLM for the target task from the VLM Zoo.
Model Selection.  Previous model selection methods typically estimate a PTM’s performance on a target task by estimating the correlation between the features extracted by the pre-trained model itself (Tran et al., 2019, 2020; You et al., 2021, 2022; Ding et al., 2022; Huang et al., 2022) or a general model (Zhang et al., 2023) and target labels on the target dataset’s samples. However, VLMs are typically used in zero-shot or few-shot scenarios, where target data is limited, making traditional model selection approaches unsuitable for VLMs. Additionally, previous methods have mainly focused on single-modal models, overlooking the characteristics of VLMs. Therefore, in this paper, we concentrate on designing model selection algorithms that are suitable for scenarios with limited data and take into account the characteristics of VLMs.

### v3

@@ 3.2 Bridge the Modality Gap and Capability Gap
The modified text embedding $\tilde{\bm{t}}_{m,i}$ serves as better image proxies. In other words, classification metrics on $f^{T}_{m}(t_{i})$ only reveal the discerning ability of the text encoder of $f_{m}$, which is far from the (cross-modal) zero-shot classification ability due to the modality gap. By bridging such a gap with modified text embedding, classification metrics on $\tilde{\bm{t}}_{m,i}$ are closer to the classification metrics on images with textual classifier. Therefore, we use $\{\tilde{\bm{t}}_{m,1},\cdots\}$ in Equation 6 as better inputs to calculate $\bm{s}_{m}^{\mathcal{T}}$ in Equation 3. The updated score vectors $\bm{s}_{m}^{\mathcal{T}}$ are then input to the ranker model $f_{R}$, which is able to get more accurate VLM’s performance prediction $\hat{p}_{m}^{\mathcal{T}}$. Based on the performance prediction $\{\hat{p}_{m}^{\mathcal{T}}\}_{m=1}^{M}$, we can obtain the VLMs’ ranking $\bm{\hat{r}}^{\mathcal{T},(1)}=[\hat{r}_{1}^{\mathcal{T},(1)},\cdots,\hat{r}_{M}^{\mathcal{T},(1)}]=\text{Ranking}\left(\left[\hat{p}_{1}^{\mathcal{T}},\cdots,p_{M}^{\mathcal{T}}\right]\right)$. $\text{Ranking}(\cdot)$ transforms the accuracies into models’ ranking.
@@
Bridge the Capability Gap. Whether the $m$-th VLM $f_{m}$ fits the target task $\mathcal{T}$ could also be estimated by the performance of $f_{m}$ on the open-source datasets related to $\mathcal{T}$. We first calculate the VLM’s class-level performance ranking over the whole open-source datasets. Given the $k$-th open-source class $c_{k}^{\mathcal{S}}$ and the corresponding set of images $D_{\mathcal{S}_{k}}^{I}$, we calculate the zero-shot classification accuracy $p_{m,k}^{\mathcal{S}}$ via Equation 1. We then determine each VLM’s ranking on the $k$-th open-source class $c_{k}^{\mathcal{S}}$ using Equation 7. We calculate VLMs’ ranking for other open-source classes in the same way.
$\bm{r}^{\mathcal{S}}_{k}=[r_{1,k}^{\mathcal{S}},\cdots,r_{M,k}^{\mathcal{S}}]=\text{Ranking}\left(\left[p_{1,k}^{\mathcal{S}},\cdots,p_{M,k}^{\mathcal{S}}\right]\right)\;.$ (7)
@@
Next, we estimate the ranking of a certain VLM $f_{m}$ on the target dataset $\mathcal{T}$. By re-organizing the ranking values in Equation 7, the performance ranking vector of $f_{m}$ on all $k_{\mathcal{S}}$ open-source classes is $\bm{r}_{m}^{\mathcal{S}}=[r_{m,1}^{\mathcal{S}},\cdots,r_{m,k_{\mathcal{S}}}^{\mathcal{S}}]\in\mathbb{Z}_{+}^{k_{\mathcal{S}}}$. If $f_{m}$ ranks high on certain open-source classes related to the classes in the target dataset, the performance ranking of $f_{m}$ on the target dataset may also be high. Thus, we perform a weighted sum of ranking values in $\bm{r}_{m}^{\mathcal{S}}$ and assign larger weights to those classes related to the target dataset. This can be done by using $\bm{\gamma}^{*}$:
$\hat{\bm{r}}_{m}^{\mathcal{T}}=\bm{r}_{m}^{\mathcal{S}}\;\bm{\gamma}^{*}\;.$ (8)
@@
Elements in $\hat{\bm{r}}_{m}^{\mathcal{T}}\in\mathbb{R}^{k_{\mathcal{T}}}$ are the predicted ranking of the m-th VLM $f_{m}$ for $k_{\mathcal{T}}$ target classes. Since we only need the relative order of ranks, there is no additional scale factor in Equation 8. After that, we average class-specific ranking values in $\hat{\bm{r}}_{m}^{\mathcal{T}}$, and use $\hat{r}_{m}^{\mathcal{T},(2)}=\text{Mean}\left(\hat{\bm{r}}_{m}^{\mathcal{T}}\right)$ to denote the estimated ranking of $f_{m}$ on the target dataset $\mathcal{T}$. In summary, we predict the ranking of models on each class in the target dataset based on their rankings in classes of the open-source datasets, guided by the semantic relevance between the classes. We predict the ranking of models on each category in the target dataset based on their rankings in categories of the open-source dataset, guided by the semantic relevance between the categories. Benefiting from the models’ consistently aligned classification performance ranking across similar classes, this approach bridges the capability gap in VLM selection.
@@ 3.3 Summary of Swab
As is described in Section 3.2, given a target dataset $\mathcal{T}$ and a VLM $f_{m}$, we denote the two performance ranking predictions obtained by bridging the modality gap and capability gap as $\hat{r}_{m}^{\mathcal{T},(1)}$ and $\hat{r}_{m}^{\mathcal{T},(2)}$, respectively. $\hat{r}_{m}^{\mathcal{T},(1)}$ is predicted based on ModelGPT with our modified embeddings of the generated auxiliary texts for $\mathcal{T}$. $\hat{r}_{m}^{\mathcal{T},(2)}$ is predicted by the weighted sum of VLM’s class-specific ranking values on the open-source datasets. These two predictions, respectively, originate from learning-based and non-learning-based LOVM methods. We ensemble two predictions together and achieve a more accurate model ranking estimation. We utilize the weighted Borda Count to aggregate two rankings:
$\hat{r}^{\mathcal{T},\rm ens}_{m}=\alpha\cdot\hat{r}_{m}^{\mathcal{T},(1)}+\left(1-\alpha\right)\cdot\hat{r}_{m}^{\mathcal{T},(2)}.$ (9)
We set $\alpha=0.5$. Ultimately, Swab determines the final predicted ranking of the VLMs in the VLM Zoo based on $\hat{r}^{\mathcal{T},\rm ens}_{m}$. The pseudo-code of Swab is listed in Algorithm 1.
@@ 4.1 Evaluation on LOVM Benchmark
Setups. We follow LOVM [64] to use its provided VLM Zoo. In addition, to further enhance the diversity of the VLM Zoo, we add some representative VLMs such as BLIP [28] and BEiT-3 [52] to expand the VLM Zoo. The final VLM Zoo contains a total of 43 models, which differ in aspects such as model architecture, pre-training datasets, and training methods. In the appendix, we also provide experimental results of different model selection methods on the 35 models originally offered by LOVM. We evaluate different methods on 23 datasets, i.e. ImageNet [8], Aircraft [34], CIFAR100 [26] and so on. We obtain VLM’s ground truth ranking based on VLM’s Top-1 Accuracy calculated on the target dataset’s test image set.
@@
Baseline. We select representative methods for each of the three paradigms mentioned in Section 2.2 as our baselines. For the first paradigm, we use four classic model selection methods: H-Score [2], NCE [48], LEEP [49] and LogME [58]. For the second paradigm, we use the VLM’s ranking on ImageNet (INB) and VLM’s average ranking (Avg Rank) on classes of the open-source datasets in the LOVM Benchmark. For the third paradigm, we compare our method with ModelGPT [64].
@@
Evaluations. We use Top-5 Recall and Kendall’s Rank Correlation to measure the similarity between the predicted and the ground truth model rankings to evaluate the LOVM method’s performance. We also calculate the sum of these two metrics to consider the method’s comprehensive capability.
@@
Implementation Details. For a fair comparison, Swab follow ModelGPT [64] to sequentially extract a target dataset from each of the 23 datasets in the LOVM Benchmark and treat the remaining datasets as open-source datasets. Besides, Swab adopts ModelGPT’s approach of adding Gaussian noise to corrupt the target dataset’s generated text embeddings. Since LOVM does not provide the specific image data for the 23 datasets, we download these datasets ourselves and adopt their standard data splits. Using the templates provided by LOVM, we construct classifiers and recalculate each VLM’s zero-shot image classification accuracy on these datasets. Additionally, we utilize the code provided by LOVM to generate class-related text data using ChatGPT. For H-Score, NCE, LEEP, LogME, INB, and Avg Rank, we follow the practices of previous work and do not add noise to the model’s inputs. To ensure reliable results, we conduct ten repeated experiments using random seeds from 1 to 10 and report the mean value and standard deviation of ModelGPT’s performance and Swab’s performance in Table 1.
Results Analysis. From Table 1, we can draw the following conclusions: (1) Metric-based non-learning model selection methods such as LogME show poor performance on the LOVM Benchmark. This is primarily because such algorithms rely on the target dataset’s images, thus the modality gap has a greater negative impact on them when using generated text data as a substitute for images. (2) Using open-source datasets is helpful for LOVM. We find that using open-source datasets in a non-learning way (e.g. INB, Avg Rank) or a learning way (e.g. ModelGPT) all helps LOVM, since their performance significantly surpasses that of methods not utilizing open-source datasets (e.g. LogME). (3) Despite leveraging more open-source datasets, the performance of Average Rank is worse than INB. This confirms our analysis of the Capability Gap, which suggests a discrepancy between the average ranking of a VLM and its ranking on a specific dataset. (4) Our Swab achieves the best performance across all evaluation metrics. Notably, our final performance of $R_{5}+\tau$ (0.822) represents a significant improvement of 14.8% over the original SoTA method ModelGPT (0.716).
@@
Table 1: Results on LOVM Benchmark. We evaluate our method across 23 datasets and 43 pre-trained VLMs. The results are averaged over all datasets. Our Swab achieves the best results across all metrics. For methods that involve adding random noise to data features, we report the standard deviation of metrics across 10 experiments to mitigate the impact of randomness on result reliability.
Methods H-Score NCE LEEP LogME INB Avg Rank ModelGPT Swab
$R_{5}(\uparrow)$ 0.174 0.235 0.161 0.191 0.443 0.443 $\text{0.446}\scriptstyle\pm\text{0.004}$ $\textbf{0.504}\scriptstyle\pm\text{0.000}$
$\tau(\uparrow)$ $0$0.014 0.014 -0.014 0.267 0.246 $\text{0.270}\scriptstyle\pm 0.009$ $\textbf{0.318}\scriptstyle\pm\text{0.002}$
$R_{5}$ + $\tau(\uparrow)$ 0.174 0.221 0.175 0.177 0.710 0.689 $\text{0.716}\scriptstyle\pm\text{0.011}$ $\textbf{0.822}\scriptstyle\pm\text{0.002}$
Table 2: Ablation Study of Swab. Swab-C, Swab-M, and Swab indicates only bridging the Capability Gap, only bridging the Modality Gap, and bridging both gaps in Swab.
@@
Swab-C $\text{0.487}\scriptstyle\pm\text{0.012}$ $\text{0.296}\scriptstyle\pm\text{0.018}$ $\text{0.783}\scriptstyle\pm\text{0.019}$
Swab-M $\text{0.474}\scriptstyle\pm\text{0.006}$ $\text{0.316}\scriptstyle\pm\text{0.019}$ $\text{0.790}\scriptstyle\pm\text{0.017}$
Swab $\textbf{0.504}\scriptstyle\pm\text{0.000}$ $\textbf{0.318}\scriptstyle\pm\text{0.002}$ $\textbf{0.822}\scriptstyle\pm\text{0.002}$
@@ 4.2 Ablation Study
We conduct an ablation study to demonstrate that bridging the Modality Gap and Capability Gap are both essential for Swab. Table 2 presents our experiment results, from which we can observe that Swab achieves the best performance across all metrics when both gaps are bridged simultaneously. The ablation study confirms our analysis.
4.3 Influence of Key Components in Swab
Will Bridge the Capability Gap Be Beneficial for VLM Selection? We compare the LOVM performance directly using the VLM’s average ranking on each class of open-source datasets and weighted-sum ranking based on transport matrix $\bm{\gamma}^{*}$. The results are shown in subsection 4.3. We can find that utilizing class relevance to bridge the Capability Gap is beneficial for VLM’s Model Selection.
Will Bridge the Modality Gap Be Beneficial for VLM Selection? To eliminate the interference of other factors, we solely utilize the learning-based predicted rankings $\hat{r}^{\mathcal{T},(1)}_{m}$ in Swab, and the input to the ranker model $f_{m}$ only consists of metrics calculated on the generated text data $\tilde{D}_{\mathcal{T}}$, which serves as substitutes for images. In this way, the method’s performance depends solely on the quality of the generated text data $\tilde{D}_{\mathcal{T}}$. From the subsection 4.3, we can find that the generated text data $\tilde{D}_{\mathcal{T}}$ become better substitutes for image data after bridging the Modality Gap.
Table 3: Results of $\hat{r}^{\mathcal{T},(2)}_{m}$ on the LOVM before and after bridging the Capability Gap.
@@ 4.3 Influence of Key Components in Swab
Average Rank 0.443 0.246 0.689
OT Weighted Rank 0.513 0.217 0.730
Table 4: Results of $\hat{r}^{\mathcal{T},(2)}_{m}$ before and after bridging the Modality Gap (MG).
@@
Before Bridging MG 0.216 0.061 0.277
After Bridging MG 0.371 0.080 0.451
Which Kind of Gap Vectors Should We Use? When utilizing the gap vectors from open-source datasets, we have two options: (1) Use the dataset-level mean gap vector calculated on the whole dataset’s image-text pairs. (2) Use the class-level mean gap vector calculated on the corresponding class’s image-text pairs. We hope that the gap vectors are as close to each other as possible so that their mean vector can substitute well for the whole set. Based on this idea, we calculate the statistics of the gap vectors within a dataset and within each class. We calculate three metrics which include: (1) the standard deviation of these gap vectors’ magnitude; (2) the mean cosine similarity between these gap vectors and their corresponding mean gap vectors; and (3) the standard deviation of these cosine similarities. These metrics reflect the consistency of the gap vectors. subsection 4.3 shows the results.
Table 5: Results of metrics measuring gap vectors’ consistency belonging to the same dataset or the same class. M: Magnitude, D: Direction.
@@
Class Mean 0.03 0.85 0.04
Table 6: Results of Swab-M on the LOVM Benchmark using the dataset-level mean gap vectors and class-level mean gap vectors.
@@
Dataset Mean 0.443 0.304 0.747
Class Mean 0.474 0.316 0.790
From the subsection 4.3, we can find that the class-level gap vectors tend to be more consistent, which inspires us to use the class-level mean gap vectors. We also compare the results of Swab-M on the LOVM Benchmark using the dataset-level mean gap vectors and the class-level mean gap vectors, respectively. The implementation details are the same as Table 2. subsection 4.3 shows the results, which verifies that using the class-level mean gap vectors is a better choice.